Find the minimum and maximum wavelengths of sound in water that is in the audible range
(20–20000 Hz) for an average human ear. Speed of sound in water 1450 m/s.
 

Grade:12

1 Answers

Arun
6 years ago
for minimum wavelength n = 20 KHz ⇒v = nλ⇒ λ = (1450//20 * 10^3) = 7.25 cm. b) for maximum wavelength n should be minimum ⇒ v = nλ⇒ λ = v/n ⇒ 1450 / 20 = 72.5 m.